How many moles are in 3.90 g of (NH4)2SO4?
Answer in units of mol.
1
Expert's answer
2018-11-15T04:54:49-0500

Answer on Question #82900, Chemistry/ General Chemistry

How many moles are in 3.90 g of (NH4)2SO4? Answer in units of mol.

Solution

Molar mass (NH4)2SO4 : 2(14.01) + 8(1.01) + 32.07 + 4(16.00) = 132.17 g (NH4)2SO4/ mol (NH4)2SO4

n = m/M

n((NH4)2SO4) = 3.90 g /132.17 g/mol = 0.0295 mol

Answer: 0.0295 mol